OpenStudy (anonymous):

(2x^2 + 6x) / (10x^3)

OpenStudy (anonymous):

x+3 over 5x^2

OpenStudy (ash2326):

@kymy_rose00 did you understand?

OpenStudy (anonymous):

is that all the work?

OpenStudy (anonymous):

\[x+3\div5x ^{2}\]

OpenStudy (anonymous):

how did you guys get that?

OpenStudy (ash2326):

\[\frac{(2x^2 + 6x) } { (10x^3)}\] To simplify this we need to cancel the common factors from the numerator and the denominator We see that 2x is common in both terms of numerator \[\frac{(2x\times x + 3\times 2x.) } { (10x^3)}\] and 2x is also a factor of the denominator \[\frac{(2x\times x + 3\times 2x.) } { (2x \times 5x^2)}\] Now you cancel 2x from both numerator and denominator!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ash your so smart :)

OpenStudy (ash2326):

Thanks @Bambi and welcome @kymy_rose00